# ReDit: Reward Dithering for Improved LLM Policy Optimization

### v2

@@ Appendix B Theorems and proofs
[…]
@@
B.3 Proof of Proposition 6.3
The proof of Proposition 6.3 is expressed as follows:
Proof.
Suppose the original reward model is $r_{RM}(\mathbf{x},\mathbf{y})$. We add a zero-mean Gaussian random variable $\epsilon$ with variance $\sigma^{2}$ to it, resulting in the new reward signal $\tilde{r}_{RM}(\mathbf{x},\mathbf{y})$:
$\tilde{r}_{RM}(\mathbf{x},\mathbf{y})=r_{RM}(\mathbf{x},\mathbf{y})+\epsilon,$ (19)
where $\epsilon\sim\mathcal{N}(0,\sigma^{2})$, and we assume that $\epsilon$ is independent of $\mathbf{y}$ (conditioned on $\mathbf{x}$).
1. Change in Reward Variance (Based on Definition 2).
Definition 2 defines the reward variance induced by the reward model $r_{RM}$ under policy $\pi_{\theta}$ and prompt $\mathbf{x}$ as:
$\text{Var}_{\mathbf{y}\sim\pi_{\theta}(\cdot|\mathbf{x})}[r_{RM}(\mathbf{x},\mathbf{y})]:=\mathbb{E}_{\mathbf{y}\sim\pi_{\theta}(\cdot|\mathbf{x})}\left[\left(r_{RM}(\mathbf{x},\mathbf{y})-\mathbb{E}_{\mathbf{y}^{\prime}\sim\pi_{\theta}(\cdot|\mathbf{x})}[r_{RM}(\mathbf{x},\mathbf{y}^{\prime})]\right)^{2}\right].$ (20)
This is the standard definition of variance: the expected squared deviation from the mean.
Now, we compute the variance of the perturbed reward signal $\tilde{r}_{RM}$:
$\text{Var}_{\mathbf{y}\sim\pi_{\theta}(\cdot|\mathbf{x})}[\tilde{r}_{RM}(\mathbf{x},\mathbf{y})]=\text{Var}_{\mathbf{y}\sim\pi_{\theta}(\cdot|\mathbf{x})}[r_{RM}(\mathbf{x},\mathbf{y})+\epsilon].$ (21)
Using the property of variance for two random variables $A$ and $B$:
$\text{Var}(A+B)=\text{Var}(A)+\text{Var}(B)+2\text{Cov}(A,B).$ (22)
Here, $A=r_{RM}(\mathbf{x},\mathbf{y})$ (a random variable depending on $\mathbf{y}$), and $B=\epsilon$ (independent of $\mathbf{y}$). Since $\epsilon$ is independent of $\mathbf{y}$, $A$ and $B$ are independent, so $\text{Cov}(A,B)=0$. Therefore:
$\text{Var}_{\mathbf{y}\sim\pi_{\theta}(\cdot|\mathbf{x})}[r_{RM}(\mathbf{x},\mathbf{y})+\epsilon]=\text{Var}_{\mathbf{y}\sim\pi_{\theta}(\cdot|\mathbf{x})}[r_{RM}(\mathbf{x},\mathbf{y})]+\text{Var}(\epsilon).$ (23)
Since $\text{Var}(\epsilon)=\sigma^{2}$, the perturbed reward variance becomes:
$\text{Var}_{\mathbf{y}\sim\pi_{\theta}(\cdot|\mathbf{x})}[\tilde{r}_{RM}(\mathbf{x},\mathbf{y})]=\text{Var}_{\mathbf{y}\sim\pi_{\theta}(\cdot|\mathbf{x})}[r_{RM}(\mathbf{x},\mathbf{y})]+\sigma^{2}.$ (24)
Conclusion: If $\sigma^{2}>0$, adding zero-mean Gaussian noise to the reward model increases the reward variance by exactly $\sigma^{2}$. This aligns with our intuitive understanding based on Theorem 3.3.
2. Change in Reward Model Accuracy (Based on Definition 1).
Definition 1 defines the accuracy of a reward model $r_{RM}$ at a given prompt $\mathbf{x}$ and distribution $\mathcal{D}$ over unordered output pairs as:
$\text{acc}_{\mathbf{x},\mathcal{D}}(r_{RM}):=\mathbb{E}_{(\mathbf{y},\mathbf{y}^{\prime})\sim\mathcal{D}}\left[\mathbbm{1}\left[\operatorname{sign}(r_{RM}(\mathbf{x},\mathbf{y})-r_{RM}(\mathbf{x},\mathbf{y}^{\prime}))=\operatorname{sign}(r_{G}(\mathbf{x},\mathbf{y})-r_{G}(\mathbf{x},\mathbf{y}^{\prime}))\right]\right],$ (25)
where $r_{G}$ is the ground truth reward function. This measures the probability that the reward model correctly ranks a pair $(\mathbf{y},\mathbf{y}^{\prime})$ relative to the ground truth.
Now consider the accuracy of the perturbed reward model $\tilde{r}_{RM}$:
$\text{acc}_{\mathbf{x},\mathcal{D}}(\tilde{r}_{RM})=\mathbb{E}_{(\mathbf{y},\mathbf{y}^{\prime})\sim\mathcal{D}}\left[\mathbbm{1}\left[\operatorname{sign}(\tilde{r}_{RM}(\mathbf{x},\mathbf{y})-\tilde{r}_{RM}(\mathbf{x},\mathbf{y}^{\prime}))=\operatorname{sign}(r_{G}(\mathbf{x},\mathbf{y})-r_{G}(\mathbf{x},\mathbf{y}^{\prime}))\right]\right].$ (26)
We analyze the difference:
$\tilde{r}_{RM}(\mathbf{x},\mathbf{y})-\tilde{r}_{RM}(\mathbf{x},\mathbf{y}^{\prime})=(r_{RM}(\mathbf{x},\mathbf{y})+\epsilon_{1})-(r_{RM}(\mathbf{x},\mathbf{y}^{\prime})+\epsilon_{2}),$ (27)
where $\epsilon_{1},\epsilon_{2}\sim\mathcal{N}(0,\sigma^{2})$ are independently sampled. Let:
$\Delta r_{RM}=r_{RM}(\mathbf{x},\mathbf{y})-r_{RM}(\mathbf{x},\mathbf{y}^{\prime}),\quad\Delta r_{G}=r_{G}(\mathbf{x},\mathbf{y})-r_{G}(\mathbf{x},\mathbf{y}^{\prime}),$ (28)
and define $\eta=\epsilon_{1}-\epsilon_{2}\sim\mathcal{N}(0,2\sigma^{2})$. Then:
$\tilde{r}_{RM}(\mathbf{x},\mathbf{y})-\tilde{r}_{RM}(\mathbf{x},\mathbf{y}^{\prime})=\Delta r_{RM}+\eta.$ (29)
The condition for accuracy becomes:
$\operatorname{sign}(\Delta r_{RM}+\eta)=\operatorname{sign}(\Delta r_{G}),$ (30)
compared to the original condition:
$\operatorname{sign}(\Delta r_{RM})=\operatorname{sign}(\Delta r_{G}).$ (31)
We now analyze how the addition of noise affects this condition:
- If the original model is accurate, i.e., $\operatorname{sign}(\Delta r_{RM})=\operatorname{sign}(\Delta r_{G})$, then adding noise $\eta$ may cause $\operatorname{sign}(\Delta r_{RM}+\eta)\neq\operatorname{sign}(\Delta r_{G})$, especially when $|\eta|$ is large enough to flip the sign of $\Delta r_{RM}$.
- If the original model is inaccurate, i.e., $\operatorname{sign}(\Delta r_{RM})\neq\operatorname{sign}(\Delta r_{G})$, there is a small chance that $\eta$ flips the sign back to match $\operatorname{sign}(\Delta r_{G})$, but this is not systematic and depends on the values of $\Delta r_{RM}$ and $\Delta r_{G}$.
In general, since $\eta$ is independent of both $\Delta r_{RM}$ and $\Delta r_{G}$, it is more likely to disrupt the correct sign relationship rather than correct it. Thus, adding random noise tends to make the ranking decisions more random.
Conclusion: Adding zero-mean Gaussian noise to the reward model makes its relative ranking of output pairs more random, thereby reducing the reward model’s accuracy (acc).
3. Upper and lower bounds of training time).
According to Theorem 3.3, the lower bound on the training time $t_{\gamma}$ is influenced by the variance of the reward function. Specifically, using Equation (24), we derive that the lower bound on $t_{\gamma}$ satisfies:
$\Omega\left(\left(\mathbb{E}_{q\sim p_{Q}}\left[\text{Var}_{o\sim\pi_{\theta}(\cdot|q)}R(q,o)\right]+\sigma^{2}\right)^{-\frac{1}{3}}\right)\leq t_{\gamma}.$ (32)
According to Theorem 3.4, in some cases (particularly when a perfectly accurate model would lead to arbitrarily slow training), an inaccurate model could instead lead to faster increases in the true reward. We derive that the lower bound on $t_{\gamma}$ satisfies:
$t_{\gamma}\leq\mathcal{O}(\pi_{\theta(0)}(\mathbf{y}^{\gamma}|\mathbf{x})^{-1})$ (33)
@@ D.3 Specific experimental parameters
Appendix E More result
@@ E.1 Main Result
Tables  7,  8,  9 show the comparison of ReDit on different datasets. ReDit significantly improves the convergence speed of GRPO. At any same step, ReDit achieves better performance.
E.2 Baseline Result
@@
E.3 Different LLMs Result
In this section, we present all numerical results in Fig. 6.
Table 11: Llama-3.2-3B-Instruct performance comparison of different training steps on the GSM8K dataset
[…]
Instruct model 73.62 - - - - - - - - -
GRPO - 73.60 73.34 74.74 74.27 73.90 73.90 73.90 74.01 74.04
Uniform ReDit - 74.81 74.81 74.89 75.05 75.05 75.66 76.13 76.27 76.27
Gauss ReDit - 74.30 74.52 75.13 75.13 75.13 76.48 76.40 76.25 76.25
Table 12: Llama-3.1-8B-Instruct performance Comparison of Different Training Steps on the GSM8K Dataset
[…]
Instruct model 81.05 - - - - - - - - -
GRPO - 81.02 81.30 81.92 81.79 82.01 82.36 82.36 81.63 82.12
Uniform ReDit - 83.81 83.24 83.45 84.32 84.01 84.32 84.12 84.73 84.92
Gauss ReDit - 83.80 83.44 83.92 84.47 84.86 84.62 84.12 84.35 84.12
Table 13: Mistral-7B-Instruct-v0.3 performance Comparison of Different Training Steps on the GSM8K Dataset
[…]
Instruct model 53.68 - - - - - - - - -
GRPO - 56.33 58.00 57.70 57.39 58.23 58.23 57.92 58.91 59.14
Uniform ReDit - 58.25 58.48 58.45 59.83 59.68 59.23 60.29 61.45 62.07
Gauss ReDit - 58.57 58.85 58.68 59.68 59.29 60.68 60.61 61.61 61.76
Table 14: Ministral-8B-Instruct-2410 performance Comparison of Different Training Steps on the Geometry3K Dataset
[…]
Instruct model 82.34 - - - - - - - - -
GRPO - 82.94 83.85 84.53 84.84 84.84 85.22 84.61 84.15 83.70
Uniform ReDit - 85.39 85.78 85.69 85.69 85.91 86.90 86.90 87.01 87.23
Gauss ReDit - 85.24 85.46 85.44 85.44 86.22 86.44 86.69 86.76 86.12
E.4 Reward Model Result
In this section, we present all numerical results in Fig. 10.
Table 15: Performance Comparison at Different Training Steps on Reward Model
[…]
Instruct model 84.91 - - - - - - - - -
RM - 85.03 86.01 86.73 87.06 87.53 87.56 87.87 88.03 88.21
RM + Gauss ReDit - 85.13 85.99 85.79 86.37 87.36 87.36 87.24 87.96 88.12
RM + Uniform ReDit - 85.76 85.63 87.03 86.32 87.13 87.53 88.01 88.63 88.63
E.5 Direct Gradient Manipulation Result
In this section, we present all numerical results in Fig. 11.
Table 16: Performance Comparison at Different Training Steps on Different Gradient Manipulation
[…]
Instruct model 84.91 - - - - - - - - -
GRPO - 85.03 86.01 86.73 87.06 87.53 87.56 87.87 88.03 88.21
Gradient Clipping - 85.65 86.32 86.34 86.37 87.12 87.75 88.31 88.57 88.57
Dynamic Sampling - 85.15 86.12 86.12 86.43 87.01 87.03 87.34 87.92 88.01
Uniform ReDit - 89.16 89.16 89.31 89.31 89.31 89.99 89.99 89.99 90.76
Gauss ReDit - 89.02 89.37 89.61 89.54 89.54 89.54 89.61 89.61 90.46
E.6 Variance Result
In this section, we show more results on the performance of ReDit as the perturbation changes. As shown in Figure  15, the variance of uniform perturbation is similar to the variance of Gaussian perturbation, and the appropriate variance can achieve the best performance. The specific numerical results are shown in Tables  17 and  18.
@@ E.6 Variance Result
E.7 Scheduled Perturbation Result

### v4

@@ B.3 Proof of Proposition 2
$\text{Var(noise gradient)}=\text{Var}\left(\epsilon\cdot\nabla_{\theta}\log\pi_{\theta}\right)=\sigma^{2}\cdot\mathbb{E}\left[\|\nabla_{\theta}\log\pi_{\theta}(o|q)\|^{2}\right]>0,$ (19)
@@ D.3 Specific experimental parameters
Appendix E Additional experiments
E.1 Results on the Code Generation Datasets
To validate the general applicability of our method (ReDit) beyond mathematical reasoning, we conducted a comprehensive evaluation on the domain of code generation. We performed experiments on three widely-used coding benchmarks: APPS, HumanEval, and CodeContests. This evaluation was designed to test the hypothesis that the core benefit of ReDit—stabilizing the learning signal to improve optimization—is a general principle that is not limited to a single domain.
The results are presented in 15. The plots consistently demonstrate that both the Gaussian (Gauss 0.05) and Uniform (Uniform 0.05) variants of ReDit significantly and consistently outperform the GRPO baseline across all three coding benchmarks. On all datasets, our method not only achieves a higher final $pass@1$ accuracy but also exhibits a faster convergence rate. This strongly suggests that ReDit provides a more robust optimization pathway, and its benefits generalize effectively to complex tasks such as code generation.
Figure 15: Performance comparison on three code generation benchmarks: (left) APPS test, (center) HumanEval test, and (right) CodeContests test. The $pass@1$ accuracy is reported across training steps. Both ReDit variants (Gauss 0.05 and Uniform 0.05) consistently and significantly outperform the GRPO baseline, confirming the general applicability of our method to the coding domain.
E.2 Results on Full Parameter Fine-Tuning
To confirm that the benefits of ReDit are not limited to parameter-efficient fine-tuning (PEFT) methods like LoRA, we conducted additional experiments using a full parameter fine-tuning approach. This evaluation addresses whether ReDit’s effectiveness is a general property of the optimization process itself, rather than an artifact of a specific tuning method [56].
For these experiments, we utilized a setup that differs from our primary PEFT experiments; specifically, we employed the VERL framework for training with 8 GPUs. We evaluated this full fine-tuning setup on our three mathematical reasoning benchmarks: GSM8K, MATH, and Geo3k.
The results are presented in 16. The plots clearly demonstrate that ReDit (both Gauss 0.05 and Uniform 0.05 variants) consistently outperforms the GRPO baseline across all three benchmarks in this demanding full-tuning setting. The performance gap is particularly notable on the MATH and Geo3k datasets, where the GRPO baseline shows signs of stagnation, while our method continues to improve. These findings confirm that ReDit is a robust and general-purpose technique, delivering consistent performance gains in both parameter-efficient and full fine-tuning paradigms.
Figure 16: Performance comparison on mathematical reasoning benchmarks using full parameter fine-tuning. The plots show test accuracy across training steps for (left) GSM8K, (center) MATH, and (right) Geo3k. In this full-tuning setting, both ReDit variants (Gauss 0.05 and Uniform 0.05) consistently achieve higher test accuracy than the GRPO baseline, confirming that our method’s benefits generalize beyond parameter-efficient tuning (PEFT).
E.3 Results on DeepSeek Distillation Models
A potential concern regarding our primary results is that they are predominantly focused on the Qwen2.5 model family. To further demonstrate the robustness and architectural generalizability of ReDit, we conducted additional experiments to validate its effectiveness on specialized reasoning models.
Specifically, we evaluated our method on two expert distillation models: DeepSeek-R1-Distill-Llama-8B and DeepSeek-R1-Distill-Qwen-7B. We used the GSM8K benchmark to compare their mathematical reasoning performance against the GRPO baseline. The results are presented in 17. The plots clearly show that both the Gaussian (Gauss 0.05) and Uniform (Uniform 0.05) variants of ReDit consistently outperform the GRPO baseline on both DeepSeek models.
This finding is significant as it confirms that ReDit’s ability to stabilize the training signal and improve performance is a general principle. It is not limited to a single model family but holds true across various model architectures, including those specifically optimized for reasoning tasks.
Figure 17: Evaluating ReDit’s generalizability on specialized reasoning models. The plots show Test Accuracy (%) on the GSM8K benchmark for (left) DeepSeek-R1-Distill-Llama-8B and (right) DeepSeek-R1-Distill-Qwen-7B. These results confirm that ReDit’s performance advantage over the GRPO baseline holds across different model architectures, not just the Qwen models used in the main experiments.
Appendix F More result
@@ F.1 Main Result
Tables  7,  89 show the comparison of ReDit on different datasets. ReDit significantly improves the convergence speed of GRPO. At any same step, ReDit achieves better performance.
F.2 Baseline Result
@@
F.3 Variance Result
In this section, we show more results on the performance of ReDit as the perturbation changes. As shown in Figure  18, the variance of uniform perturbation is similar to the variance of Gaussian perturbation, and the appropriate variance can achieve the best performance. The specific numerical results are shown in Tables  11 and  12.
@@ F.3 Variance Result
F.4 Scheduled Perturbation Result